# Evaluating the Diversity and Quality of LLM Generated Content

### v1

@@ 4 Experimental Setup
Let $\mathcal{T}=\{t_{1},\ldots,t_{L}\}$ be a fixed set of test inputs. For any generated program $g$, let $g(t_{l})$ denotes its execution on test case $t_{l}\in\mathcal{T}$, and $o_{l}$ denotes the corresponding output. We define the validity function $V:\mathcal{G}\to\{0,1\}$ as follows:
$V(g)=\begin{cases}1&\text{if }g(t_{l})\text{ executes without error and produces non-null output }\forall l\in\{1,\dots,L\},\\ 0&\text{otherwise.}\end{cases}$
@@
For the remaining diversity metrics, we adopt Equation 3 and specify the semantic distance function $d_{\text{sem}}(g,g^{\prime})$ used. For Lexical diversity, we use Expectation-Adjusted Distinct $n$-grams (EAD) with $n$-grams of length $n$$=$4. For syntactic diversity, we adapt the Distinct-N metric to the Abstract Syntax Tree (AST) of each generated program: to isolate the syntactic structure of a program (e.g., for-loop instead of recursion) from superficial choices (e.g., variable names), we canonicalize all program identifiers. For two programs, we calculate the ratio of the number of unique subtrees of height $H$ across both programs to the total number of subtrees of height $H$ in both programs, where $H$$=$4. In order to measure neural diversity, we adapt existing methods of neural diversity metrics (Tevet & Berant, 2021) to our domain using CodeBertScore (Zhou et al., 2023). We use CodeBertScore because it closely resembles the models used in the NLP literature to evaluate diversity (Tevet & Berant, 2021) and has been evaluated as effective in the evaluation of neural codes. We also attempted to use CodeLLama-7B-Instruct embeddings and ICEScore (Zhuo, 2024) to evaluate diversity and found CodeLLama similar to CodeBertScore and ICEScore to correlate strongly with temperature, such as the lexical diversity metric. See the the additional results in Section A.1.
@@
Models and Sampling. The main empirical questions we seek to answer in our work surround the effects of different post-training algorithms and model size on diversity (SFT, DPO, PPO, GRPO, etc). In order to isolate the effects of different post-training strategies, we utilize the Tulu-2 (Ivison et al., 2023) and Llama-2 (Touvron et al., 2023) as well as the Tulu-3 (Lambert et al., 2024) and Llama3.1 (Dubey et al., 2024) families of models. Tulu-2 consists of post-trained Llama-2 models, and Tulu-3 consists of post-trained Llama3.1 models. In Table 1, we provide a categorization of the post-training methods. Using these we can compare a SFT-tuned model to a DPO-tuned model from the same family. For each problem description $x_{i}$ in our dataset, we generate $K=$ 32 programs, yielding 3,456 total programs sampled for each experiment. In cases where models are post-trained with numerous algorithms, we often categorize the model with the most aggressive algorithm used, e.g. we would categorize a model adapted with DPO and PPO as PPO.
@@
Prompt selection. The choice of prompt can affect the nature of generations. Because of this, we created three separate prompt templates: a zero-shot prompt, a two-shot prompt, and a two-shot prompt with chain-of-thought reasoning. This design allows us to probe generation behavior across a variety of settings. The few-shot examples included in the prompts were simple, manually written examples shared across all problems in the dataset. We provide the examples in Section A.5.
@@ 5.1 Effect of Post-Training and Preference-Tuning on Diversity
Table 2: Model Comparison Results. Results from Wilcoxon’s Signed-Rank Test p-values: W (p), and Effect Size measured by Cohen’s D: ES (d). Bold p-values are below 0.05, and bold d-values have an absolute value greater than 0.8 (large effect size). For effect sizes, positive values indicate the second model type in the comparison has higher values. "Inst" = Instruction-tuned (Post-Trained), "Pref" = Preference-tuned and combines RL and DPO models. The columns correspond to metrics in Section 4 (e.g. “Semantic" corresponds to effective semantic diversity).
@@
Table 3: Validity-Controlled Model Comparison Results. Bold p-values are below 0.05, and bold d-values have an absolute value greater than 0.8 (large effect size)
Preference tuning: more semantic diversity at the cost of less diversity in form. In Table 2, we summarize our results across all diversity metrics when comparing base models and their instruction-tuned counterparts, as well as comparisons between post-training techniques. Each row is a comparison between two different fine-tuning approaches; for each metric, the columns show the statistical significance W (p) and effect size ES (d). A statistically significant, positive effect means that the second post-training strategy is greater than the first in terms of that metric.
We find a very clear trend: all post-training techniques increase both effective semantic diversity and validity over base models. We also find that RL in particular improves dramatically on SFT when it comes to effective semantic diversity. Additionally, we find an interesting pattern where preference-tuning broadly induces a dramatic decrease in the syntactic and lexical diversity of generations.
Semantic diversity driven by higher quality. In Table 3, we perform similar statistical tests as in Table 2, only with the subset of generations that were valid. We find that within this subset preference-tuned models generally have more semantic duplicates than their SFT-tuned counterparts. This suggests that while preference-tuning, especially RL, is liable to increase semantic duplicates in the subset of high quality generations, the volume of high quality generations more than compensates for this effect. Lastly, while we do not find a statistically-significant pattern between DPO and RL for effective semantic diversity, we find strong evidence and large effect sizes that RL-tuned models have less syntactic and lexical diversity than associated DPO-tuned models.
@@ 5.2 Effects of Model Size
Larger models increase semantic diversity without reducing the diversity of form. In Table 2, we show the results when comparing small and large models of the same family. We generally see higher semantic diversity in larger models without sacrificing lexical/syntactic diversity. We believe there is an intuitive justification: larger models have a higher capacity to model a greater amount of distributions in its corpus and their relative strengths in helpfulness and quality is well-documented.
Smaller models may be more compute-efficient to generate a unique example. While we find that larger models may be more diverse per-generation, we explore the question of which models are most efficient with regard to a computation budget. Consider a researcher trying to generate a large number of unique, synthetic programs: is it more advantageous to take fewer samples from a very-large LLM, many samples from a smaller LM, or is there a “sweet spot" in-between? In Figure 4 we explore this question by plotting the parameter-efficiency for diversity vs. the model’s size, where this efficiency is the effective semantic diversity using Equation 2 of a model normalized by the number of parameters of a model. For our fixed sampling budget of 32 generations, smaller models are consistently more efficient on a parameter basis. We believe that further work should be done to investigate if this pattern continues to hold as the number of samples increases to ensure semantically unique programs do not saturate faster for certain model classes than others. This is one of many potential directions for future work that could further leverage our framework.
Figure 4: Model efficiency: We plot the parameter efficiency of a model in generating unique examples vs. model size (log-scale).
@@ 6 Discussion and Conclusion
We have proposed a novel strategy for studying effective semantic diversity where we leverage code execution to measure and balance quality and diversity. Using this methodology, we perform extensive empirical analysis of LLM diversity. We find counterintuitive insights into the nature of how factors like post-training and model size meaningfully affect diversity. We encourage further work broadening empirical questions asked and exploring domains like natural language in the future.
@@ A.1 Additional Neural Diversity Metrics
In addition to using CodeBertScore, we also demonstrate neural diversity with respect to temperature sweeps similar to those in Figure 1 using CodeLlama-7B-Instruct (Roziere et al., 2023) embeddings. We also use an LLM-as-a-judge to calculate pairwise similarity using ICEScore (Zhuo, 2024) using gpt-4o-mini as the LLM-judge. We find that ICEScore correlates heavily with temperature and does not seem to reflect quality. Similarly with CodeLlama embeddings, the Llama8B-base model’s generations are consistently considered more diverse than the Instruct variants despite the base model generally producing far more low-quality and invalid outputs.
@@ A.2 Sample Size Confounding Diversity and Analysis of Pairwise Diversity Metric
In practice, we found that varying the size of the subset under consideration could dramatically affect the calculation of Equation 2: for the same model, as samples tended to get larger, this metric would decrease. When we use this metric for most of our experiments, it is sufficient, because we keep the number of samples constant. However, when we analyze the subset of valid-only programs, the size of this population may vary from model to model depending on the average quality of generations. Thus it is necessary to use Equation 3.
Analysis of Pairwise Diversity Metric. We provide an analysis of Equation 3. For a given large language model $f$, we assume that only a finite number $K$ of distinct semantic meanings can be generated by $f$. We first establish that the original semantic diversity metric converges to zero as the number of sampled responses tends to infinity. Specifically, the original semantic diversity metric is defined as
@@
Next, we show that the new metric defined in Equation 3, converges to a constant as $n\to\infty$. As before, we assume that there are $K$ distinct semantic meanings in total, and let $\pi_{k}$ denote the proportion of responses corresponding to the $k$-th semantic meaning. This implies that the number of times each semantic meaning is sampled is $\pi_{k}n$, where $\sum_{k=1}^{K}\pi_{k}=1$. Thus, we have
@@ A.3 Dataset Creation and Additional Details
In our experiments, we instruct LLMs to return their outputs. We maintain a wrapper script that executes the function $f(...)$ for each input, obtains the returned result, and then serializes the object as well as its type (recursively) into a string. We then utilize these strings as the basis for our semantic comparisons. For each generation, we execute all test cases and capture the resulting outputs.
@@
Because LLMs often generate natural language to accompany generated programs, extracting programs from the generations is a non-trivial task, especially for pre-trained models. We developed an extraction utility that extracts not only the target function $f(...)$, but also any helper functions and imports that may be relevant. To safely execute programs at scale, we perform all execution inside an isolated Docker container to prevent adverse consequences of blindly executing LLM outputs.

### v2

@@ 4 Experimental Setup
Let $\mathcal{T}=\{t_{1},\ldots,t_{L}\}$ be a fixed set of test inputs. For any generated program $g$, let $g(t_{l})$ denotes its execution on test case $t_{l}\in\mathcal{T}$, and $o_{l}$ denotes the corresponding output. Specifically, for code we define the validity function $V:\mathcal{G}\to\{0,1\}$ as follows:
[…]
@@
In the domain of natural language, we prompt a LLM judge with a set of criteria to judge the quality of the creative generation and then normalize the score to fall between $[0,1]$.
@@
For natural language, for $Sim$, the semantic similarity function that does not explicitly consider quality, we prompt an LLM judge to evaluate the conceptual and thematic overlap between two generations and normalize the score. For all LLM judge tasks we utilize gpt-4.1-mini-2025-04-14. Due to the scale of our experiments, we subsampled 32 pairs with replacement from all possible pairs for each set of generations.
For the remaining diversity metrics, we adopt Equation˜3 and utilize the following distance functions in place of $d_{\text{sem}}(g,g^{\prime})$. For Lexical diversity, we use Expectation-Adjusted Distinct $n$-grams (EAD) with $n$-grams of length nn==4. For syntactic diversity in code, we adapt the Distinct-N metric to the Abstract Syntax Tree (AST) of each generated program: to isolate the syntactic structure of a program (e.g., for-loop instead of recursion) from superficial choices (e.g., variable names), we canonicalize all program identifiers. For two programs, we calculate the ratio of the number of unique subtrees of height $H$ across both programs to the total number of subtrees of height $H$ in both programs, where HH==4. In order to measure neural diversity for code, we adapt existing methods of neural diversity metrics (Tevet & Berant, 2021) to our domain using CodeBertScore (Zhou et al., 2023). We use CodeBertScore because it closely resembles the models used in the NLP literature to evaluate diversity (Tevet & Berant, 2021) and has been evaluated as effective in the evaluation of neural codes. We also attempted to use CodeLLama-7B-Instruct embeddings and ICEScore (Zhuo, 2024) to evaluate diversity and found CodeLLama similar to CodeBertScore and ICEScore to correlate strongly with temperature, such as the lexical diversity metric. See the the additional results in Section˜A.1. For natural language, we report the average score from the LLM judge $\mathcal{D}$ before normalizing it by the generation’s quality.
@@
Models and Sampling. The main empirical questions we seek to answer in our work surround the effects of different post-training algorithms and model size on diversity (SFT, DPO, PPO, GRPO, etc). In order to isolate the effects of different post-training strategies, we utilize the Tulu-2 (Ivison et al., 2023) and Llama-2 (Touvron et al., 2023) as well as the Tulu-3 (Lambert et al., 2024) and Llama3.1 (Dubey et al., 2024) families of models. Tulu-2 consists of post-trained Llama-2 models, and Tulu-3 consists of post-trained Llama3.1 models. In Table˜1, we provide a categorization of the post-training methods. Using these we can compare a SFT-tuned model to a DPO-tuned model from the same family. For each problem description $x_{i}$ in our dataset, we generate $K=$ 32 programs, yielding 3,456 total programs sampled for each experiment. In cases where models are post-trained with numerous algorithms, we often categorize the model with the most aggressive algorithm used, e.g. we would categorize a model adapted with DPO and PPO as PPO.
@@
Prompt selection. The choice of prompt can affect the nature of generations. Because of this, we created three separate prompt templates: a zero-shot prompt, a two-shot prompt, and a two-shot prompt with chain-of-thought reasoning. This design allows us to probe generation behavior across a variety of settings. The few-shot examples included in the prompts were simple, manually written examples shared across all problems in the dataset. We provide the examples in Section˜A.6.
@@ 5.1 Effect of Post-Training and Preference-Tuning on Diversity
Table 2: Model Comparison Results. Results from Wilcoxon’s Signed-Rank Test p-values: W (p), and Effect Size measured by Cohen’s D: ES (d). Bold p-values are below 0.05, and bold d-values have an absolute value greater than 0.8 (large effect size). For effect sizes, positive values indicate the second model type in the comparison has higher values. "Inst" = Instruction-tuned (Post-Trained), "Pref" = Preference-tuned and combines RL and DPO models. The columns correspond to metrics in Section˜4 (e.g. “Semantic" corresponds to effective semantic diversity).
Validity Semantic Lexical Neural
Comparison W (p) ES (d) W (p) ES (d) W (p) ES (d) W (p) ES (d)
Base vs. Inst (All) <0.001 2.02 <0.001 1.50 <0.001 1.65 <0.001 -3.60
Base vs. Inst-SFT <0.001 2.19 0.001 1.66 <0.001 2.53 <0.001 -2.61
Base vs. Inst-DPO 0.001 1.51 0.014 0.89 <0.001 2.27 <0.001 -4.27
Base vs. Inst-RL <0.001 4.88 <0.001 3.91 0.022 0.73 <0.001 -5.68
Base vs. Inst-Pref <0.001 2.18 <0.001 1.53 <0.001 1.34 <0.001 -4.58
SFT vs. DPO <0.001 1.52 0.001 0.96 <0.001 1.12 <0.001 -1.61
SFT vs. RL 0.004 5.79 0.008 2.76 0.004 3.60 0.004 -4.67
SFT vs. Pref <0.001 2.36 <0.001 1.54 <0.001 1.48 <0.001 -2.39
DPO vs. RL 0.055 0.51 0.025 0.59 0.82 0.28 0.02 -0.40
Sm. vs. Lg. <0.001 0.41 <0.001 0.53 0.04 -0.21 0.009 -0.20
Sm. vs. Lg. Base 0.031 3.04 0.031 3.35 0.844 0.05 0.062 -1.21
Sm. vs. Lg. Inst 0.001 0.40 0.006 0.42 0.003 -0.32 0.055 -0.26
Table 3: Creative Writing (Natural Language) Model Comparison Results. Bold p-values are below 0.05, and bold d-values have an absolute value greater than 0.8 (large effect size)
@@
Table 4: Validity-Controlled Model Comparison Results for Code. Bold p-values are below 0.05, and bold d-values have an absolute value greater than 0.8 (large effect size)
Higher effecitve semantic diversity introduced by preference tuning. We summarize results across all diversity metrics when comparing base models and their instruction-tuned counterparts, as well as when comparing different post-training techniques, for programming tasks in Table˜2 and natural language tasks in Table˜3. Each row compares two fine-tuning approaches; for each metric, the columns report the statistical significance W (p) and effect size ES (d). A statistically significant positive effect indicates that the second post-training strategy outperforms the first for that metric.
We observe a clear trend: all post-training techniques increase both effective semantic diversity and validity relative to base models. RL methods, in particular, yield substantial improvements over SFT in effective semantic diversity. We also find an interesting pattern: preference tuning tends to substantially reduce syntactic and lexical diversity in programming tasks, yet increases these metrics in natural language creative writing. These results suggest that in domains requiring high-quality and diverse outputs, preference-tuned models can outperform both SFT and base models. Furthermore, in creative writing—where diverse word choice and stylistic variety are often desirable—preference-tuned models may hold an advantage in stylistic capabilities. Additional experiments are provided in Appendix A.7 and Appendix A.8.
Semantic diversity driven by higher quality. In Table˜4, we conduct statistical tests similar to those in Table˜2 for code, but restrict the analysis to only valid generations. We find that within this subset, preference-tuned models generally have more semantic duplicates than their SFT-tuned counterparts. This suggests that while preference tuning, particularly RL, tends to increase semantic duplication within high-quality generations, the overall increase in the number of high-quality generations more than compensates for this effect.
@@ 5.2 Effects of Model Size
Larger models increase semantic diversity. In Table˜2 and Table˜3, we compare small and large models within the same family. In both code and natural language, larger models generally exhibit higher semantic diversity. In code, this increase does not come at the expense of lexical or syntactic diversity; however, in natural language, we observe a small but significant decrease in lexical diversity. We attribute the improvements in quality and effective semantic diversity to the well-documented strengths of larger models in helpfulness and quality.
Smaller models may be more compute-efficient for generating unique examples. While larger models tend to produce more diverse outputs per generation, we also examine which models are most efficient given a fixed computational budget. Consider a researcher aiming to generate a large number of unique synthetic programs: is it better to take fewer samples from a very large LLM, many samples from a smaller LM, or find a “sweet spot" in between? In Figure˜3, we explore this question by plotting the parameter efficiency for diversity against model size, where efficiency is measured as the effective semantic diversity (per Equation˜2) normalized by the number of model parameters. For a fixed sampling budget of 32 generations per prompt in the programming dataset, smaller models are consistently more efficient on a per-parameter basis. We believe that further work should be done to investigate whether this pattern continues to hold as the number of samples increases, to ensure semantically unique programs do not saturate faster for certain model classes than others.
Figure 3: Model efficiency: We plot the parameter efficiency of a model in generating unique examples vs. model size (log-scale) on our programming dataset.
@@ 6 Discussion and Conclusion
We propose a novel strategy for studying effective semantic diversity by leveraging code execution to jointly measure and balance quality and diversity. In the natural language domain, we use LLM judges as a proxy for human evaluation. Using these methodologies, we conduct an extensive empirical analysis of LLM diversity and find counterintuitive insights into how factors such as post-training and model size meaningfully influence diversity. We encourage further work to further broaden the range of empirical questions explored in this area.
Acknowledgements
We would like to thank Christopher Watson, Cassidy Laidlaw, and other members of Berkely AI Research for feedback on our draft. This material is partly based on research sponsored in part by the Air Force Research Laboratory (agreement number FA8750-19-2-0200 and award W911NF-20-1-0080), an Amazon Research Award Fall 2023, and an Amazon/ASSET Gift for Research in Trustworthy AI. The U.S. Govt. is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of the Air Force Research Laboratory or the U.S. Government.
@@ A.1 Additional Neural Diversity Metrics
In addition to using CodeBertScore, we also demonstrate neural diversity with respect to temperature sweeps similar to those in Figure˜1 using CodeLlama-7B-Instruct (Roziere et al., 2023) embeddings. We also use an LLM-as-a-judge to calculate pairwise similarity using ICEScore (Zhuo, 2024) using gpt-4o-mini as the LLM-judge. We find that ICEScore correlates heavily with temperature and does not seem to reflect quality. Similarly with CodeLlama embeddings, the Llama8B-base model’s generations are consistently considered more diverse than the Instruct variants despite the base model generally producing far more low-quality and invalid outputs.
@@ A.2 Sample Size Confounding Diversity and Analysis of Pairwise Diversity Metric
In practice, we found that varying the size of the subset under consideration could dramatically affect the calculation of Equation˜2: for the same model, as samples tended to get larger, this metric would decrease. When we use this metric for most of our experiments, it is sufficient, because we keep the number of samples constant. However, when we analyze the subset of valid-only programs, the size of this population may vary from model to model depending on the average quality of generations. Thus it is necessary to use Equation˜3.
Analysis of Pairwise Diversity Metric. We provide an analysis of Equation˜3. For a given large language model $f$, we assume that only a finite number $K$ of distinct semantic meanings can be generated by $f$. We first establish that the original semantic diversity metric converges to zero as the number of sampled responses tends to infinity. Specifically, the original semantic diversity metric is defined as
@@
Next, we show that the new metric defined in Equation˜3, converges to a constant as $n\to\infty$. As before, we assume that there are $K$ distinct semantic meanings in total, and let $\pi_{k}$ denote the proportion of responses corresponding to the $k$-th semantic meaning. This implies that the number of times each semantic meaning is sampled is $\pi_{k}n$, where $\sum_{k=1}^{K}\pi_{k}=1$. Thus, we have
@@ A.3 Dataset Creation and Additional Details
In our experiments, we instruct LLMs to return their outputs. We maintain a wrapper script that executes the function f()f(...) for each input, obtains the returned result, and then serializes the object as well as its type (recursively) into a string. We then utilize these strings as the basis for our semantic comparisons. For each generation, we execute all test cases and capture the resulting outputs.
@@
Because LLMs often generate natural language to accompany generated programs, extracting programs from the generations is a non-trivial task, especially for pre-trained models. We developed an extraction utility that extracts not only the target function f()f(...), but also any helper functions and imports that may be relevant. To safely execute programs at scale, we perform all execution inside an isolated Docker container to prevent adverse consequences of blindly executing LLM outputs.
A.4 Raw Experimental Results for All Experiments
In Table˜5 and Table˜6, we include all our raw results used in our analysis.
Model Prompt Validity Semantic Lexical Neural
Llama-3.1-70B-Instruct Zero-Shot 69.42 11.32 97.72 29.12
Llama-3.1-70B-Instruct Two-Shot 36.28 6.02 98.56 49.98
Llama-3.1-70B-Instruct Two-Shot CoT 33.15 5.24 98.83 52.69
Llama-3.1-Tulu-3-70B Zero-Shot 91.95 15.95 98.92 21.20
Llama-3.1-Tulu-3-70B Two-Shot 90.31 18.63 98.90 26.17
Llama-3.1-Tulu-3-70B Two-Shot CoT 89.71 19.63 99.04 30.73
Llama-3.1-Tulu-3-70B-DPO Zero-Shot 90.18 17.87 99.05 24.92
Llama-3.1-Tulu-3-70B-DPO Two-Shot 90.10 17.65 99.02 24.86
Llama-3.1-Tulu-3-70B-DPO Two-Shot CoT 89.39 20.11 98.95 32.27
Llama-3.1-Tulu-3-70B-SFT Zero-Shot 68.33 22.83 98.42 56.75
Llama-3.1-Tulu-3-70B-SFT Two-Shot 58.09 18.22 98.71 57.09
Llama-3.1-Tulu-3-70B-SFT Two-Shot CoT 59.98 17.17 98.90 51.50
Llama-3.1-70B Zero-Shot 56.12 17.97 96.51 59.86
Llama-3.1-70B Two-Shot 40.72 10.07 98.13 67.81
Llama-3.1-70B Two-Shot CoT 52.35 17.76 98.07 69.17
Llama-3.1-8B-Instruct Zero-Shot 37.28 5.99 98.13 47.89
Llama-3.1-8B-Instruct Two-Shot 24.11 2.81 99.24 51.81
Llama-3.1-8B-Instruct Two-Shot CoT 24.31 3.05 99.14 53.22
Llama-3.1-Tulu-3.1-8B Zero-Shot 90.55 17.83 98.89 24.17
Llama-3.1-Tulu-3.1-8B Two-Shot 87.77 21.40 98.71 34.59
Llama-3.1-Tulu-3.1-8B Two-Shot CoT 92.37 18.56 98.87 21.84
Llama-3.1-Tulu-3-8B Zero-Shot 90.72 18.96 99.05 24.83
Llama-3.1-Tulu-3-8B Two-Shot 85.47 19.08 98.87 33.14
Llama-3.1-Tulu-3-8B Two-Shot CoT 91.11 18.49 98.97 26.63
Llama-3.1-Tulu-3-8B-DPO Zero-Shot 88.84 18.13 98.84 30.67
Llama-3.1-Tulu-3-8B-DPO Two-Shot 86.59 19.84 98.82 34.71
Llama-3.1-Tulu-3-8B-DPO Two-Shot CoT 89.83 19.51 98.94 29.30
Llama-3.1-Tulu-3-8B-SFT Zero-Shot 58.60 19.76 98.47 60.15
Llama-3.1-Tulu-3-8B-SFT Two-Shot 55.41 18.75 98.70 58.21
Llama-3.1-Tulu-3-8B-SFT Two-Shot CoT 57.70 19.55 98.81 57.43
Llama-3.1-8B Zero-Shot 32.20 9.85 96.39 69.74
Llama-3.1-8B Two-Shot 28.69 6.31 98.68 71.12
Llama-3.1-8B Two-Shot CoT 37.89 11.80 98.57 74.33
Llama-2-70b-chat-hf Zero-Shot 30.37 8.05 97.76 64.03
Llama-2-70b-chat-hf Two-Shot 23.02 3.86 99.20 63.88
Llama-2-70b-chat-hf Two-Shot CoT 24.64 4.34 99.24 65.06
tulu-2-70b Zero-Shot 75.60 17.52 98.40 51.31
tulu-2-70b Two-Shot 71.17 14.54 98.74 52.00
tulu-2-70b Two-Shot CoT 73.37 17.88 98.69 54.56
tulu-2-dpo-70b Zero-Shot 78.64 15.52 98.73 44.07
tulu-2-dpo-70b Two-Shot 75.10 16.63 98.86 47.46
tulu-2-dpo-70b Two-Shot CoT 78.60 17.68 98.84 49.00
Llama-2-70b-hf Zero-Shot 22.87 7.11 96.85 75.34
Llama-2-70b-hf Two-Shot 21.30 4.93 98.71 75.63
Llama-2-70b-hf Two-Shot CoT 29.24 8.90 98.59 77.89
Llama-2-7b-chat-hf Zero-Shot 22.39 4.98 98.34 67.20
Llama-2-7b-chat-hf Two-Shot 16.54 2.44 99.48 67.81
Llama-2-7b-chat-hf Two-Shot CoT 17.28 2.77 99.45 69.08
tulu-2-7b Zero-Shot 53.99 12.95 98.07 66.84
tulu-2-7b Two-Shot 49.64 11.01 98.56 67.74
tulu-2-7b Two-Shot CoT 51.51 12.86 98.47 69.61
tulu-2-dpo-7b Zero-Shot 64.67 13.41 98.49 58.75
tulu-2-dpo-7b Two-Shot 59.52 12.74 98.74 60.59
tulu-2-dpo-7b Two-Shot CoT 62.36 14.19 98.67 62.58
Llama-2-7b-hf Zero-Shot 19.71 4.83 97.60 77.03
Llama-2-7b-hf Two-Shot 17.70 3.44 99.08 77.63
Llama-2-7b-hf Two-Shot CoT 23.52 5.98 98.93 79.62
Table 5: Raw Results for All Natural Language Experiments
Model Prompt Validity Semantic Lexical Syntactic Neural
Llama-3.1-70B-Instruct Zero-Shot 37.36 32.96 67.19 71.98 5.60
Llama-3.1-70B-Instruct Two-Shot 11.89 11.40 66.47 65.19 4.51
Llama-3.1-70B-Instruct Two-Shot CoT 14.29 13.48 66.77 67.39 4.97
Llama-3.1-Tulu-3-70B Zero-Shot 39.55 32.41 73.91 78.09 8.88
Llama-3.1-Tulu-3-70B Two-Shot 41.49 33.22 72.76 75.33 4.43
Llama-3.1-Tulu-3-70B Two-Shot CoT 40.88 32.85 73.26 76.69 8.31
Llama-3.1-Tulu-3-70B-DPO Zero-Shot 38.57 32.93 75.37 79.81 9.54
Llama-3.1-Tulu-3-70B-DPO Two-Shot 38.27 32.83 74.45 78.27 6.78
Llama-3.1-Tulu-3-70B-DPO Two-Shot CoT 39.18 33.81 74.89 78.93 9.07
Llama-3.1-Tulu-3-70B-SFT Zero-Shot 37.36 32.96 67.19 71.98 5.60
Llama-3.1-Tulu-3-70B-SFT Two-Shot 35.48 29.37 68.76 71.38 5.23
Llama-3.1-Tulu-3-70B-SFT Two-Shot CoT 36.42 31.12 68.48 72.15 5.40
Llama-3.1-70B Zero-Shot 20.18 24.02 54.05 55.47 3.51
Llama-3.1-70B Two-Shot 13.28 14.70 57.87 57.00 3.46
Llama-3.1-70B Two-Shot CoT 15.17 18.00 57.25 58.00 3.64
Llama-3.1-8B-Instruct Zero-Shot 15.47 18.98 56.40 60.12 3.41
Llama-3.1-8B-Instruct Two-Shot 8.54 10.37 56.09 57.54 3.22
Llama-3.1-8B-Instruct Two-Shot CoT 9.94 12.70 55.63 58.68 3.35
Llama-3.1-Tulu-3.1-8B Zero-Shot 37.67 32.41 73.32 77.66 8.91
Llama-3.1-Tulu-3.1-8B Two-Shot 37.06 32.67 72.45 76.10 6.83
Llama-3.1-Tulu-3.1-8B Two-Shot CoT 38.57 33.70 73.19 77.53 8.73
Llama-3.1-Tulu-3-8B Zero-Shot 38.27 32.93 74.52 78.64 9.22
Llama-3.1-Tulu-3-8B Two-Shot 36.73 31.85 73.19 76.69 6.20
Llama-3.1-Tulu-3-8B Two-Shot CoT 37.97 32.85 73.91 77.83 8.66
Llama-3.1-Tulu-3-8B-DPO Zero-Shot 37.06 32.41 74.06 78.28 8.91
Llama-3.1-Tulu-3-8B-DPO Two-Shot 36.42 31.70 73.19 76.69 6.83
Llama-3.1-Tulu-3-8B-DPO Two-Shot CoT 37.67 32.56 73.75 77.53 8.44
Llama-3.1-Tulu-3-8B-SFT Zero-Shot 35.18 29.37 68.17 71.38 5.23
Llama-3.1-Tulu-3-8B-SFT Two-Shot 34.85 29.67 68.01 71.18 5.40
Llama-3.1-Tulu-3-8B-SFT Two-Shot CoT 35.48 30.11 68.32 71.78 5.57
Llama-3.1-8B Zero-Shot 11.89 15.99 47.23 48.41 2.78
Llama-3.1-8B Two-Shot 9.64 12.41 51.05 51.11 2.73
Llama-3.1-8B Two-Shot CoT 11.28 14.70 50.43 51.51 2.91
Llama-2-70b-chat-hf Zero-Shot 11.28 15.84 52.30 55.87 3.30
Llama-2-70b-chat-hf Two-Shot 6.43 8.52 51.82 53.57 3.11
Llama-2-70b-chat-hf Two-Shot CoT 7.82 10.67 52.14 54.36 3.25
tulu-2-70b Zero-Shot 31.86 29.52 64.74 68.35 5.06
tulu-2-70b Two-Shot 29.27 26.22 64.58 66.62 4.57
tulu-2-70b Two-Shot CoT 30.57 28.52 64.90 67.79 4.95
tulu-2-dpo-70b Zero-Shot 33.26 30.41 67.03 70.28 5.40
tulu-2-dpo-70b Two-Shot 31.25 27.70 66.71 68.35 4.79
tulu-2-dpo-70b Two-Shot CoT 32.56 29.37 66.87 69.52 5.23
Llama-2-70b-hf Zero-Shot 8.24 12.11 44.15 45.33 2.59
Llama-2-70b-hf Two-Shot 6.43 9.96 46.58 47.02 2.54
Llama-2-70b-hf Two-Shot CoT 8.24 11.26 46.27 47.42 2.68
Llama-2-7b-chat-hf Zero-Shot 8.85 12.56 49.21 52.31 3.03
Llama-2-7b-chat-hf Two-Shot 5.82 7.78 48.58 50.18 2.84
Llama-2-7b-chat-hf Two-Shot CoT 7.22 9.67 48.89 51.11 2.97
tulu-2-7b Zero-Shot 24.39 24.46 58.25 61.63 4.13
tulu-2-7b Two-Shot 22.56 21.70 58.09 59.89 3.68
tulu-2-7b Two-Shot CoT 23.78 23.41 58.41 60.89 3.98
tulu-2-dpo-7b Zero-Shot 26.83 26.07 61.08 64.36 4.46
tulu-2-dpo-7b Two-Shot 24.39 23.41 60.61 62.43 3.98
tulu-2-dpo-7b Two-Shot CoT 25.91 25.19 60.92 63.56 4.30
Llama-2-7b-hf Zero-Shot 6.73 10.22 41.72 42.70 2.38
Llama-2-7b-hf Two-Shot 5.52 8.37 43.84 44.33 2.32
Llama-2-7b-hf Two-Shot CoT 6.73 9.52 43.68 44.53 2.46
Table 6: Raw Results for General Code Experiments